OpenStudy (anonymous):

How do you reduce the following: x-y/x^2-1 times x-1/x^2-y^2?

OpenStudy (anonymous):

@satellite73 @Zarkon

OpenStudy (anonymous):

20 gabriel13 0@satellite73 @Zarkon

OpenStudy (zarkon):

this... \[\frac{x-y}{x^2-1}\frac{x-1}{x^2-y^2}\]?

OpenStudy (anonymous):

YES

OpenStudy (zarkon):

notice that \[x^2-1=(x-1)(x+1)\] and \[x^2-y^2=(x-y)(x+y)\]

OpenStudy (zarkon):

\[\frac{x-y}{x^2-1}\frac{x-1}{x^2-y^2}\] \[=\frac{x-y}{(x+1)(x-1)}\frac{x-1}{(x-y)(x+y)}\] \[=\frac{x-y}{(x+1)\cancel{(x-1)}}\frac{\cancel{x-1}}{(x-y)(x+y)}\] \[=\frac{x-y}{(x+1)}\frac{1}{(x-y)(x+y)}\] \[=\frac{\cancel{x-y}}{(x+1)}\frac{1}{\cancel{(x-y)}(x+y)}\] \[=\frac{1}{(x+1)}\frac{1}{(x+y)}=\frac{1}{(x+1)(x+y)}\]

OpenStudy (anonymous):

yes is this thanks

OpenStudy (anonymous):

a^3+a2b/5a.25/3b+3a help me @Zarkon's

OpenStudy (anonymous):

@Zarkon's help me please

OpenStudy (anonymous):

a^3+a2b/5a.25/3b+3a